Chevalier Manufacturing issued a callable bond with a 2026 maturity date. The bond was sold at par and the call premium was set equal to the coupon rate of 20%. A declining call premium was also in place so that the premium declined evenly over the last ten years to a premium of zero. What would be the call premium if the bond was called in 2017?

$ 18

$ 20

$180

$200

cannot calculate without market price
Answer #1

call premium=1000*20%=200

By 2017, the bond has passed one year of its life, so the premium will reduce by=200/10=20

What would be the call premium if the bond was called in 2017

=call premium-reduce premium for one year

=200-20

=180

the above is answer..